OpenStudy (anonymous):

integral of b/(x(1-x)) from a to p

hartnn (hartnn):

heard of partial fractions ?

hartnn (hartnn):

\(\Large \dfrac{1}{x(x-1)}= \dfrac{A}{x}+ \dfrac{B}{x-1}\)

hartnn (hartnn):

can you find A and B ?

OpenStudy (anonymous):

Oh that absolutely helps. I just needed to see that start. Thank you.

hartnn (hartnn):

welcome ^_^ you seem to be new here \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

Haha yep I just discovered it :) thanks